OpenStudy (anonymous):

How to solve for length of a kite I need to find x where 3x+2=x+10 my problem for that bottom of the kite is how to get alone our teacher told us the answer was 4 for x but im not getting 4

OpenStudy (phi):

3x+2=x+10 this is algebra. first, add -x to both sides 3x - x + 2 = x - x + 10 simplify. on the left, you have 3 x's take away 1 x on the right side you have 1 x take away 1 x what do you get after simplifying ?